OpenStudy (anonymous):

sharyn needs 3 3/8 yards of fabric to make a dress for herself and 2 1/8 yards of the same fabric to make a matching dress for her daughter. if she buys 6 1/8 yards of fabric, how much will she have left over?

OpenStudy (anonymous):

Evaluate the following:\[\left(6+\frac{1}{8}\right)-\left(3+\frac{3}{8}\right)-\left(2+\frac{1}{8}\right) \]